OpenStudy (anonymous):

In ΔABC, if the length of side b is 3 centimeters and the measures of ∠B and ∠C are 45° and 60°, respectively, what is the length of side c to two decimal places?

OpenStudy (goformit100):

Welcome to Open Study .................. Sir/Ma'am @tarik96

OpenStudy (anonymous):

glad to be here need assistance and Im a young man

OpenStudy (campbell_st):

well you can use the law of sines \[\frac{3}{\sin(45)} = \frac{c}{\sin(60)}\]

OpenStudy (anonymous):

3.67 centimeters